A chartered jet can travel the 3600 km to a math convention in 3 h when flying with the wind. Against the wind, the trip takes 4 h. What is the jet's rate in still air? What is the rate of the wind?
Answer by jorel1380(3719)

3600/3=1200 mph
3600/4=900 mph
1200+900/2=1050 mph speed in still wind--
1200-1050=1050-900=150 mph speed of the wind..